Use the quadratic formula to determine the x-intercepts (if any) of the following function. Then evaluate the function for several values of x, and use the resulting points to graph the function.
f(x) = -x2 + 2x - 2

Answer by nyc_function(2741)

Let y = f(x)
To find the x-intercepts, let y = 0 and solve for x.
To evaluate the function, select different value for x, plug them into your function and do the math. You can use any number. I will give you 4 to play with. Let x = -2, -1, 0, 1.
To graph the function, go to mathway.com and in the graph category, write the following: Graph y = -x^2 + 2x - 2.
Is this clear?